### v1

@@ 4.6 Training and Inference
where L1 loss [49] $\mathcal{L}_{L1}$ and GIoU loss [60] $\mathcal{L}_{GIoU}$ are applied to both hand and object bounding box regression, and focal loss [61] $\mathcal{L}_{ac}$ is for interaction classification. Notably, the Ego-HOI detection task considers hand classification, which differs from the third-person perspective. Therefore, the cross-entropy loss is employed not only for object classification $\mathcal{L}_{oc}$, but also for hand classification $\mathcal{L}_{hc}$. $\lambda_{L1}$, $\lambda_{GIoU}$, $\lambda_{hoc}$ and $\lambda_{ac}$ are the hyper-parameters for weighting each loss. The loss functions of the baseline models [54, 56] and comparison models [46, 52] are similar to Eq. 10, but the details may be different due to the unique characteristics of each model.
@@
[…]
@@
And we only select the top several predictions with confidence scores above a threshold from all $N$ results.
Table 3: Performance and efficiency comparison of different HOI baselines with and without integration with our method. For clarity, all AP and Accuracy metrics are presented as percentages.
Method Backbone Ego-HOIBench Efficiency
$\mathrm{mAP}_{50}\uparrow$ $\mathrm{mAP}_{75}\uparrow$ Full $\uparrow$ Top@G Acc. $\uparrow$ #Params $\downarrow$ FPS $\uparrow$
MUREN [54] ResNet-50 83.5 75.6 65.0 81.4 75M 15.9
+ Ours ResNet-50 84.1 76.3 66.8(+1.8) 85.7(+4.3) 79M(+4M) 15.0(-5.7%)
QPIC [56] ResNet-101 84.8 80.8 76.7 86.5 60M 18.6
+ Ours ResNet-101 85.9 83.4 78.4(+1.7) 92.7(+6.2) 65M(+5M) 17.1(-8.1%)
@@ 5.1 Integration to Off-the-shelf HOI Detectors
MUREN [54] is an end-to-end Transformer-based approach with a three-branch architecture. It decouples human detection, object detection, and interaction classification, using independent decoder layers to extract task-specific tokens for sub-task learning. In our integration, the interaction branch’s attention fusion module output is leveraged as the interaction representations $\mathbf{I}$, while the human branch’s attention fusion module output serves as the hand representations $\mathbf{H}$.
QPIC [56] is one of the pioneering Transformer-based set prediction models for HOI detection. It employs a single decoder to predict all three elements of HOI: human, verb, and object. In our integration, the unified features output by the decoder are used indiscriminately as the original interaction features $\mathbf{I}$ and hand features $\mathbf{H}$. We apply a vanilla encoder to the unified features to derive the object-specific features.
@@ 5.2 Experimental Setup
Implementation Details. Our experiments cover two baselines and their integrations with our method. We also include other existing HOI detection methods for comparison, all of which are modified and retrained for the Ego-HOI detection task. Our experimental and analytical endeavors focus on the instance-level setting, as this level provides richer details. Image-level detection can be achieved by simply modifying the prediction heads or post-processing, so we do not compare them here. To obtain better detection performance, we fine-tune the object detector (usually DETR [50] with a ResNet-50 backbone) on the Ego-HOIBench training set. All experiments are performed on 4 RTX 4090 GPUs. The hyper-parameters in the experiment remain consistent with the default settings of respective methods, but the batch size and initial learning rate are adjusted according to the supported computing resources. Specifically, all experiments of HOTR [52] and MUREN [54] adopt a batch size of 16 and an initial learning rate of 5e-5. For STIP [46], the HOI detector with a frozen object detector uses a batch size of 16 and an initial learning rate of 5e-5, while the batch size is 8 and the initial learning rate is 1e-5 during the two-stage joint fine-tuning. QPIC [56] is trained with a batch size of 8 and an initial learning rate of 5e-5.
Evaluation Metrics. We evaluate models’ performance on the Ego-HOIBench benchmark using mean average precision (mAP) with IoU thresholds ranging from 0.5 to 0.95 with a step size of 0.05. A detection result is considered a true positive only if the predicted hand, object, and verb categories are all correct, and the hand and object bounding boxes have IoUs with ground truths larger than the specified threshold. We further divide all the Ego-HOI triplet categories into rare and non-rare according to whether they appear at least 100 times in the training set. Based on this criterion, we report the mAP for the Full, Rare, and Non-rare categories. The mAPs of the full testing set at IoU thresholds of 0.5 and 0.75 are reported separately, denoted as $\mathrm{mAP_{50}}$ and $\mathrm{mAP_{75}}$, similar to [44, 54]. In addition, to highlight the improvement of our method in interaction recognition, we introduce Top@G Verb Accuracy as a metric. For an image to be considered correct, the G predictions with the highest probabilities must completely cover the set of true verb labels, where G represents the number of true labels.
@@ 5.3 Improvement on Two Different Baselines
Table 3 shows the performance comparison of two mainstream baseline HOI detection methods before and after integrating our proposed method. By incorporating our approach, both baseline methods achieve significant performance improvements. Specifically, MUREN [54] achieves a 1.8% improvement in Full mAP and a 4.3% increase in Top@G Accuracy. As for QPIC [56], Full mAP is improved by 1.7% and Top@G Accuracy obtains a substantial improvement of 6.2%, setting a new high for the state-of-the-art results. These results demonstrate that our method is applicable not only to models with a unified decoder but also to the methods that decouple the sub-tasks. Moreover, our scheme imposes no specific restrictions on the backbone. Note that after integrating our module, these two baseline methods can still maintain end-to-end training and reasoning. We also compare their model sizes and runtime efficiencies to prove that the performance improvement is not due to the increase in model size. Although our method adds several million parameters, this increase is very limited relative to the original model size. Furthermore, in terms of Frames Per Second (FPS), the runtime speed drop is negligible, only a few percentage points. The results show that our technology is extremely lightweight and efficient.
Table 4: Performance comparison of our proposed method (last row) and state-of-the-art methods on the Ego-HOIBench dataset. All metrics are presented as percentages. $\dagger$ denotes that the object detector and HOI detector are further fine-tuned jointly.
@@
$\mathrm{mAP}_{50}$ $\mathrm{mAP}_{75}$ Full Rare Non-Rare Top@G Accuracy
HOTR [52] ResNet-50 44.3 18.9 22.4 9.8 27.3 85.1
STIP [46] ResNet-50 $\times$ 47.7 22.4 25.0 10.8 30.4 76.0
@@
MUREN [54] ResNet-50 83.5 75.6 65.0 61.6 66.3 81.4
QPIC [56] ResNet-50 84.4 78.7 69.5 67.6 70.3 87.4
QPIC [56] ResNet-101 84.8 80.8 76.7 75.5 77.1 86.5
QPIC + Ours ResNet-101 85.9 83.4 78.4 79.8 77.9 92.7
@@ 5.4 Comparison with State-of-the-art Methods
Table 4 presents a detailed performance comparison between our proposed method and many typical approaches, including the one-stage single-branch method QPIC [56], the one-stage two-branch method HOTR [52], the one-stage three-branch method MUREN [54], the two-stage method STIP [46] and its jointly fine-tuned version. Here, all AP and Accuracy metrics are presented in percentage form. We use QPIC as the baseline and integrate it with our scheme for comparison. Our method (last row) surpasses all existing one-stage and two-stage methods, whether in Ego-HOI detection or interaction recognition. A noteworthy phenomenon is that the rare triplet categories consistently underperform compared to the non-rare categories in terms of mAP across all other methods. In contrast, our method significantly enhances the detection performance of rare categories, even surpassing that of non-rare categories. The superior performance of our method is mainly due to the fact that we effectively extract and incorporate hand pose cues into the interaction embedding. This enhancement significantly boosts the model’s ability to distinguish complex and rare-seen interactions, further improving the overall performance of Ego-HOI detection.
@@ 5.5 Ablation Study
Table 5: Ablation study of each component in our HGIR scheme, starting from the baseline and progressively building up to our complete method. $\checkmark$ means that the corresponding component is used. HPE: Hand Pose Estimation. IR: Interactivity Refinement. HGE: Hand Geometry Extraction.
HPE IR PGE $\mathrm{mAP}_{50}$ Full Top@G Acc.
- - - 83.5 65.0 81.4
$\checkmark$ - - 83.4 66.0 81.3
$\checkmark$ $\checkmark$ - 83.6 66.1 84.7
$\checkmark$ $\checkmark$ $\checkmark$ 84.1 66.8 85.7
@@
Figure 6: Qualitative comparison between the baseline and our proposed method. For each image, the detection outputs of our proposed method are marked in green, while the baseline outputs are marked in red. The predicted classes and scores are presented in the captions. If no true positive is predicted, the score is marked as none. For clarity, the hand category is omitted from the image captions.
@@
Table 6: Performance comparison of different hand pose estimation schemes.
Method $\mathrm{mAP}_{50}$ Full Top@G Acc.
(a) Direct 82.9 65.7 85.3
(b) Learnable Point 83.1 66.6 85.3
(c) Center Point 83.6 66.5 84.3
(d) Top-Center Point 84.1 66.8 85.7
@@ 5.6 Qualitative Results and Discussions
We also compare our proposed method with the baseline across different object occlusion ratios. For Ego-HOI detection, we count prediction results according to their ground truth occlusion ratios, as shown in Fig. 7 (top). For interaction recognition, we statistically classify each prediction according to the average ground truth occlusion ratio of the instances within each image, as shown in Fig. 7 (bottom). Overall, the performance of both the baseline and our method shows a downward trend with the increase of occlusion ratio. This phenomenon occurs because occlusions may obscure critical features, thus hindering the model’s learning. Nonetheless, our method consistently outperforms the baseline at all occlusion levels. In particular, at the high occlusion level (0.8$\sim$1), our method improves Full mAP by 5.7% and Top@G Accuracy by 4.0% compared to the baseline. These significant improvements are mainly due to our method’s ability to leverage poses as additional cues to enhance interaction features and infer interactions more effectively, even when the visible portion of an object is too limited to provide enough information.
Figure 7: Performance comparison of our proposed method with the baseline for Ego-HOI detection (top) and interaction recognition (bottom) under different object occlusion ratios.
@@ 6 Conclusion
In this paper, we introduce a new benchmark dataset, Ego-HOIBench, which aims to advance the research of human-object interaction detection from an egocentric perspective. To address the detection challenges exacerbated by severe occlusions in egocentric vision, we propose a Hand Geometric and Interactivity Refinement scheme. Our method ingeniously extracts detailed hand geometric features and utilizes pose offsets prompts to refine interaction features. By fusing these complementary features, it generates more informative and contextual representations for interaction recognition. Importantly, our module can be easily integrated into existing HOI detection frameworks, making it a general and versatile solution. Extensive experiments on the Ego-HOIBench dataset demonstrate the effectiveness of our proposed approach. It significantly improves the accuracy and robustness of interaction recognition and detection compared to various baseline methods. We believe the insights from this work will inspire future research to further enhance real-world applications involving human-object interactions.
Acknowledgements

### v2

@@ 4.6 Training and Inference
where L1 loss (Ren et al., 2016) $\mathcal{L}_{L1}$ and GIoU loss (Rezatofighi et al., 2019) $\mathcal{L}_{GIoU}$ are applied to both hand and object bounding box regression. To address the class imbalance problem in interaction classification, focal loss (Lin et al., 2017) $\mathcal{L}_{ac}$ is adopted. Notably, the Ego-HOI detection task requires explicit hand classification, which differs from the third-person perspective. Therefore, the cross-entropy loss is employed for both object classification $\mathcal{L}_{oc}$ and hand classification $\mathcal{L}_{hc}$. $\lambda_{L1}$, $\lambda_{GIoU}$, $\lambda_{hoc}$ and $\lambda_{ac}$ are hyper-parameters for weighting each loss. The loss functions of the baseline models (Tamura et al., 2021; Kim et al., 2023) and comparison models (Kim et al., 2021; Zhang et al., 2022d) are similar to Eq. 11, but the details may differ due to the unique characteristics of each model.
@@
Inference. Given a set of Ego-HOI predictions $\left\{\left(\mathbf{\hat{p}}_{i}^{\mathrm{i}},\mathbf{\hat{p}}_{i}^{\mathrm{h}},\mathbf{\hat{p}}_{i}^{\mathrm{o}},\mathbf{\hat{b}}_{i}^{\mathrm{h}},\mathbf{\hat{b}}_{i}^{\mathrm{o}}\right)\right\}_{i=1}^{N}$, where $\mathbf{\hat{p}}_{i}^{\mathrm{i}}\in\mathbb{R}^{\left|\mathcal{V}\right|}$ and $\mathbf{\hat{p}}_{i}^{\mathrm{o}}\in\mathbb{R}^{\left|\mathcal{O}\right|+1}$ correspond to the classification probabilities for interaction and object respectively, the predicted category $c_{i}^{\tau}$ and its score $s_{i}^{\tau}$ are given by $\underset{k}{\arg\max\,}\mathbf{\hat{p}}_{i,k}^{\tau}$ and $\underset{k}{\max\,}\mathbf{\hat{p}}_{i,k}^{\tau}$. Considering the hand classification, the confidence score of an Ego-HOI prediction is defined as follows:
@@
We retain only the top predictions whose confidence scores exceed a predefined threshold from all $N$ predictions. Furthermore, to reduce false positives, we leverage the verb-object co-occurrence relationships shown in Fig. 3 to filter out implausible triplet combinations. This strategy ensures that the final Ego-HOI predictions are both confident and consistent with the realistic interaction patterns observed in the dataset.
@@ 5.1 Integration to Off-the-shelf HOI Detectors
MUREN (Kim et al., 2023) is an end-to-end Transformer-based approach with a three-branch architecture. It decouples human detection, object detection, and interaction classification, using independent decoder layers to extract task-specific tokens for sub-task learning. In our integration, the interaction branch’s attention fusion module output is leveraged as the interaction representations $\mathbf{I}$, while the human branch’s attention fusion module output serves as the hand representations $\mathbf{H}$.
QPIC (Tamura et al., 2021) is one of the pioneering Transformer-based set prediction models for HOI detection. It employs a single decoder to predict all three elements of HOI: human, verb, and object. In our integration, the unified features output by the decoder are used indiscriminately as the original interaction features $\mathbf{I}$ and hand features $\mathbf{H}$. We apply a vanilla encoder to the unified features to derive the object-specific features.
Table 3: Performance and efficiency comparison of different representative HOI detection baselines before and after integrating our proposed HGIR scheme on the Ego-HOIBench dataset. All AP and Accuracy metrics are reported as percentages.
Method Backbone Ego-HOIBench Efficiency
Full $\uparrow$ $\mathrm{mAP}_{50}\uparrow$ $\mathrm{mAP}_{75}\uparrow$ Top@G Acc. $\uparrow$ #Params $\downarrow$ FPS $\uparrow$
MUREN (Kim et al., 2023) ResNet-50 65.0 83.5 75.6 81.4 75M 15.9
+ HGIR (Ours) ResNet-50 66.8(+1.8) 84.1 76.3 85.7(+4.3) 79M(+4M) 15.0(-5.7%)
QPIC (Tamura et al., 2021) ResNet-101 76.7 84.8 80.8 86.5 60M 18.6
+ HGIR (Ours) ResNet-101 78.4(+1.7) 85.9 83.4 92.7(+6.2) 65M(+5M) 17.1(-8.1%)
@@ 5.2 Experimental Setup
Implementation Details. Our experiments cover two baselines and their integrations with our method. We also include other existing HOI detection methods for comparison, all of which are modified and retrained for the Ego-HOI detection task. To obtain better detection performance, we fine-tune the object detector (usually DETR (Carion et al., 2020) with a ResNet-50 backbone) on the Ego-HOIBench training set. All experiments are performed on 4 RTX 4090 GPUs. The hyper-parameters in the experiment remain consistent with the default settings of respective methods, but the batch size and initial learning rate are adjusted according to the supported computing resources. Specifically, all experiments of HOTR (Kim et al., 2021) and MUREN (Kim et al., 2023) adopt a batch size of 16 and an initial learning rate of 5e-5. For STIP (Zhang et al., 2022d), the HOI detector with a frozen object detector uses a batch size of 16 and an initial learning rate of 5e-5, while the batch size is 8 and the initial learning rate is 1e-5 during the two-stage joint fine-tuning. QPIC (Tamura et al., 2021) is trained with a batch size of 8 and an initial learning rate of 5e-5.
Evaluation Metrics. We assess the models’ performance on the Ego-HOIBench benchmark using mean average precision (mAP) with IoU thresholds ranging from 0.5 to 0.95 with a step size of 0.05. A detection result is considered a true positive only if the predicted hand, object, and verb categories are all correct, and the hand and object bounding boxes have IoUs with ground truths larger than the specified threshold. We further divide all the Ego-HOI triplet categories into rare and non-rare according to whether they appear at least 100 times in the training set. Based on this criterion, we report the mAP for the Full, Rare, and Non-rare categories. The mAPs of the full testing set at IoU thresholds of 0.5 and 0.75 are reported separately, denoted as $\mathrm{mAP_{50}}$ and $\mathrm{mAP_{75}}$, similar to (Zhang et al., 2023; Kim et al., 2023). In addition, to highlight the improvement of our method in interaction recognition, we introduce Top@G Verb Accuracy as a metric. For an image to be considered correct, the G predictions with the highest probabilities must completely cover the set of true verb labels, where G represents the number of true labels.
@@ 5.3 Improvement on Two Different Baselines
Table 3 shows the performance comparison of two mainstream baseline HOI detection methods before and after integrating our proposed method. By incorporating our approach, both baseline methods achieve significant performance improvements. Specifically, MUREN (Kim et al., 2023) achieves a 1.8% improvement in Full mAP and a 4.3% increase in Top@G Accuracy. As for QPIC (Tamura et al., 2021), Full mAP is improved by 1.7% and Top@G Accuracy obtains a substantial improvement of 6.2%, setting a new high for the state-of-the-art results. These results demonstrate that our method is applicable not only to models with a unified decoder but also to the methods that decouple the sub-tasks. Moreover, our scheme imposes no specific restrictions on the backbone. Note that after integrating our module, these two baseline methods can still maintain end-to-end training and reasoning. We also compare their model sizes and runtime efficiencies to prove that the performance improvement is not due to the increase in model size. Although our method adds several million parameters, this increase is very limited relative to the original model size. Furthermore, in terms of Frames Per Second (FPS), the runtime speed drop is negligible, only a few percentage points. The results show that our technology is extremely lightweight and efficient.
Table 4: Detailed performance comparison of our proposed method (last row) and representative state-of-the-art HOI detection methods on the Ego-HOIBench dataset. Ours adopts QPIC as a backbone followed by our proposed HGIR. All AP and Accuracy metrics are reported as percentages. $\dagger$ indicates joint fine-tuning of the object detector and HOI detector.
@@
Full Rare Non-Rare $\mathrm{mAP}_{50}$ $\mathrm{mAP}_{75}$ Top@G Accuracy
HOTR (Kim et al., 2021) ResNet-50 22.4 9.8 27.3 44.3 18.9 85.1
STIP (Zhang et al., 2022d) ResNet-50 $\times$ 25.0 10.8 30.4 47.7 22.4 76.0
@@
MUREN (Kim et al., 2023) ResNet-50 65.0 61.6 66.3 83.5 75.6 81.4
QPIC (Tamura et al., 2021) ResNet-50 69.5 67.6 70.3 84.4 78.7 87.4
QPIC (Tamura et al., 2021) ResNet-101 76.7 75.5 77.1 84.8 80.8 86.5
Ours ResNet-101 78.4 79.8 77.9 85.9 83.4 92.7
@@ 5.4 Comparison with State-of-the-art Methods
Table 4 presents a detailed performance comparison between our proposed method and many typical approaches, including the one-stage single-branch method QPIC (Tamura et al., 2021), the one-stage two-branch method HOTR (Kim et al., 2021), the one-stage three-branch method MUREN (Kim et al., 2023), the two-stage method STIP (Zhang et al., 2022d) and its jointly fine-tuned version. Here, all AP and Accuracy metrics are presented in percentage form. We use QPIC as the baseline and integrate it with our scheme for comparison. Our method (last row) surpasses all existing one-stage and two-stage methods, whether in Ego-HOI detection or interaction recognition. A noteworthy phenomenon is that the rare triplet categories consistently underperform compared to the non-rare categories in terms of mAP across all other methods. In contrast, our method significantly enhances the detection performance of rare categories, even surpassing that of non-rare categories. The superior performance of our method is mainly due to the fact that we effectively extract and incorporate hand pose cues into the interaction embedding. This enhancement significantly boosts the model’s ability to distinguish complex and rare-seen interactions, further improving the overall performance of Ego-HOI detection.
@@ 5.5 Ablation Study
Table 5: Ablation study evaluating the contribution of each component in our HGIR scheme. Starting from the baseline, components are added incrementally: Hand Pose Estimation (HPE), Interactivity Refinement (IR), and Hand Geometry Extraction (HGE). $\checkmark$ indicates the inclusion of the corresponding component.
HPE IR HGE Full $\mathrm{mAP}_{50}$ Top@G Acc.
- - - 65.0 83.5 81.4
$\checkmark$ - - 66.0 83.4 81.3
$\checkmark$ $\checkmark$ - 66.1 83.6 84.7
$\checkmark$ $\checkmark$ $\checkmark$ 66.8 84.1 85.7
@@
Table 6: Ablation study comparing different hand pose estimation schemes. “Direct” predicts joint positions directly; “Learnable Point”, “Center Point”, and “Top-Center Point” use different reference points for offset-based estimation.
Method Full $\mathrm{mAP}_{50}$ Top@G Acc.
(a) Direct 65.7 82.9 85.3
(b) Learnable Point 66.6 83.1 85.3
(c) Center Point 66.5 83.6 84.3
(d) Top-Center Point 66.8 84.1 85.7
@@
Figure 6: Qualitative comparison between the baseline and our proposed method on challenging Ego-HOI detection cases. For each image, predictions from our method are shown in green, and baseline predictions in red. The predicted triplet classes and confidence scores are shown below each image. If the model fails to output a true positive, the score is marked as none. For clarity, hand categories are omitted from the captions.
@@ 5.6 Qualitative Results and Discussions
Figure 7: Comparison of our method and the baseline on Ego-HOI detection (top) and interaction recognition (bottom) across different object occlusion ratios.
[…]
@@ 6 Conclusion
In this paper, we introduce a new benchmark dataset, Ego-HOIBench, which aims to advance the research of human-object interaction detection from an egocentric perspective. Ego-HOIBench provides high-quality, fine-grained annotations of hand-verb-object triplets, covering a diverse range of scenarios, object types, and hand configurations. To address the unique challenges of Ego-HOI detection, particularly severe occlusions and complex hand-object interactions, we propose the Hand Geometry and Interactivity Refinement (HGIR) scheme. HGIR leverages global hand pose cues and pose-aware interaction features to enhance interaction representation in a lightweight and plug-and-play manner. Extensive experiments demonstrate that our method can be seamlessly integrated into various HOI detection frameworks, consistently improving both detection and interaction recognition performance on Ego-HOIBench.
Future work will focus on addressing the imbalanced issue and further improving model generalization to rare and complex Ego-HOI triplets. We hope Ego-HOIBench and our proposed method will facilitate further research and practical advancements in egocentric vision and human-object interaction understanding.
Acknowledgments